OpenStudy (anonymous):

(x+1)^2(x+1)^4 what would it be simplifty to would it be (x+1)^7

OpenStudy (anonymous):

\[(x+1)^{2}(x+1)^{4} = (x+1)^{2+4} =(x+1)^{6}\]

OpenStudy (anonymous):

thanks